OpenStudy (anonymous):

Insert grouping symbols so that the given expression has a value of 22. 10 – 5 • 8 – 6 + 12

OpenStudy (texaschic101):

(10 - 5) * (8 - 6) + 12 = 22 5 * 2 + 12 = 22 10 + 12 = 22 22 = 22 (correct)

OpenStudy (anonymous):

:D thank you

OpenStudy (texaschic101):

no problem :)